If complex number Z^4=1-i then find all the rrots of Z.
…...................................
…...................................
…........................................
…...........................
 

Grade:11

1 Answers

Arun
6 years ago
Dear Prasad
 
z^4 = sqrt(2) (cos (- pi/4)+ i sin (- pi/4))
z^4 = sqrt(2) (cos (2k pi + (- pi/4))+ i sin (2k pi + (- pi/4)))
 
put k = 0, 1, 2 ,3
I hope that you can proceed now
 
Regards
Arun (askIITians forum expert)